OpenStudy (anonymous):

List the roots of f(x) = 24x3 – 10x2 -7x + 2. Please help me!

OpenStudy (anonymous):

Can you help me?

OpenStudy (anonymous):

Do you mean \[24x^3 + 10x^2 -7x + 2\]

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

{x = 1/4}, {x = -1/2}, {x = 2/3} use wolfram alpha